# Statistical testing of random number generators and their improvement using randomness extraction

### v1

@@ VI Conclusion and Future Work
In this work, we have presented both our $\mathsf{STE}$, a statistical testing environment to analyse the output of RNGs, and a variety of extraction methods to post-process RNGs outputs. Our $\mathsf{STE}$ software, documentation and a build file can be found at https://github.com/CQCL/random_test, and the randomness extractor software library Cryptomite can be found at https://github.com/CQCL/cryptomite and cryptomite . Our objective is to make both statistical testing and randomness extraction easy to use, efficient and openly accessible.
The results from our statistical testing tell a mixed story. First, using our $\mathsf{STE}$ we intensely tested the output of three RNGs, showing failure for two of them (reproducing and strengthening previous results hurley2020quantum ; hamburg2012analysis ). For the RNGs that failed, we observed that all of our post-processing methods improved their statistical properties, and in particular, observed that for post-processing levels level 2 or above (seeded, two-source and physical extraction), the processed output was found to be statistically indistinguishable from uniform. Unfortunately, because of the limitations of statistical testing we were unable to find non-artificial examples of RNGs failing when post-processing of level 2 (seeded extraction) was applied, but that are successful when level 3 or higher levels are applied (two-source and physical) – although level 3 is strictly stronger than level 2. The full list of our observations can be found in Section V.3 and following sections.
We could have gone even further in our numerical testing but, because numerical tests consume substantial computational resources, decided to omit certain test suites from our analysis, including SPRNG mascagni2000algorithm and Crypt-X gustafson1994computer . Moreover, we were recently made aware of the numerical tests BitReps uokbitreps and RaBiGeTe rabi , which are also not included in $\mathsf{STE}$. It would be interesting to include these in the analysis to obtain an even more intense statistical testing environment.
@@
Furthermore, it would be interesting to perform statistical testing of other RNGs with our test environment to analyse how they perform when tested beyond what is required by standardisation bodies. Similarly, it would be interesting to include different post-processing methods than the ones we presented. One could use, for example, vetted conditioning components from NIST turan2018recommendation and compare their results to the ones we obtained using information-theoretic randomness extractors.
VII Acknowledgements
We thank Erik Woodhead and Ela Lee for useful discussions and suggestions.
@@ A.1 Linear Feedback Shift Register (LFSR)
The LFSR is a class of pseudo RNG that is commonly used in applications, due to its speed and ease of implementation in both software and hardware, e.g. [60] [61]. Notably, LFSRs are used in cryptography, including in hashing and authentication [49] and stream ciphers [62]. For this work, we implement the maximal period LFSR found in [63].
This LFSR generates pseudo randomness as follows. Let $s=b_{1},\ldots,b_{32}$ denote the initial 32 bit state where $b_{i}$ denotes the $i=1,\ldots,32$th bit.
@@
    Initialize LFSR with the 32-bit initial state $s=b_{1},b_{2},...,b_{31},b_{32}$.
@@
    Repeat step (2-5) until the desired amount of bits has been generated.
@@ A.2 Intel RDSEED
Intel manufacture a hardware RNG based on thermal noise, which is present in their computer processing units. This true-RNG is constructed as follows, although a more in-depth description can be found in [10].
@@
    Initial weak randomness is generated from an entropy source. This source is a self-clocking circuit designed such that, when the clock is running, the circuit enters a meta-stable state, which then resolves to one of two possible states - determined randomly by thermal noise. The state in which the circuit resolves is the random bit output from the entropy source. This self-clocking occurs irregularly at around 3 GHz.
@@
    Health and swellness checks, which are very simple statistical tests, with the goal of detecting critical failure in the entropy source.
@@
    Processing of randomness through a cryptographic hash function.
A user then calls randomness from the true-RNG by using RDSEED. An in-depth description of all the above steps can be found in [11] and [10].
Using the Intel true-RNG, we are unable to output raw randomness from the entropy source. The best we can do is use RDSEED. In this case, some post-processing has already been performed (as described above). Some independent analysis of the quality of the Intel true-RNG has been done, including [10] and [32], where the former find that the min-entropy rate of RDSEED is around $0.65$, which is similar to our result (see Section IV).
@@ A.3 IDQ Quantis QRNG
The IDQ Quantis (USB) is QRNG based on photons hitting a 50:50 beam splitter and being detected in position 0 (reflected) and 1 (transmitted). In principle if all components are accurately modelled and the device is shielded from any outside influence, the output is perfect random numbers due to the laws of quantum mechanics. We revert the reader to the IDQ Quantis QRNG brochure for further details of its construction [22].
@@ B.1 Full results: Statistical testing
Table 9: The number of failed tests for the raw output from the 32-bit LFSR. Note, only 127/920 PractRand tests were run due to the numerous failings in the $2^{25}$ byte case. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parenthesis.
@@
Table 10: The number of failed tests for the raw output from RDSEED. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parenthesis.
@@
Table 11: The number of failed tests for the raw output from IDQ Quantis. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parenthesis.
B.2 Full results: Min entropy estimators
@@
NIST
Min-Entropy
Estimator
(/byte)
NIST
Min-Entropy
Estimator
(/bit)
@@
NIST
Min-Entropy
Estimator
(/byte)
NIST
Min-Entropy
Estimator
(/bit)
@@
Table 13: Observed NIST min-entropy estimators for RDSEED raw output.
@@
NIST
Min-Entropy
Estimator
(/byte)
NIST
Min-Entropy
Estimator
(/bit)
@@ B.2.1 Deriving a min-entropy lower bound
In this subsection, we derive a min-entropy lower bound from the observed NIST min-entropy estimators in Table 4. We use subscripts to index a single min-entropy estimator test and superscript to index the RNG which the variable refers to. Let $\mathsf{est}_{i}$ denote the $i$th observed NIST min-entropy estimate per bit for a test, $i=1,\ldots,10$, and $\overline{\mathsf{est}}$ the average estimate per bit. The sample standard deviation $\sigma$, (using Bessel’s correction), given by
@@
We compute the lower bound for min-entropy rate, $\alpha^{\mathsf{RNG}}$, including a finite statistics correction term to lower bound the true estimated min-entropy rate of each RNG with high probability. Specifically, we want
@@
where $\mathrm{\overline{est}^{\mathsf{RNG}}}$ is the average NIST min-entropy estimator for the RNG and $\sigma^{\mathsf{RNG}}$ is the observed sample standard deviation satisfies Equation 19, giving $\epsilon_{\mathsf{est}}\approx 2^{-39}$. Here, we have made the assumption that the NIST min-entropy estimator results are normally distributed (which we believe is reasonable due to each test sample being generated a significant time apart) and used the standard probability density function for normally distributed variables.
@@ C.1 Full results
Table 15: The number of failed tests for the output of the 32-bit LFSR after post-processing with the Von Neumann extractor. Note, only 127/920 PractRand tests were run due to the numerous failings in the $2^{25}$ byte case. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parenthesis.
@@
Table 16: The number of failed tests for the output of RDSEED after post-processing with the Von Neumann extractor. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parenthesis.
@@
Table 17: The number of failed tests for the output of IDQ Quantis after post-processing with the Von Neumann extractor. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parenthesis.
@@ D.1 Full results
Table 18: Statistical test results for the 32-bit LFSR as the weak input source to the strong seeded randomness Circulant extractor. The seed is randomness generated from the NIST Randomness Beacon. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parenthesis.
@@
Table 19: Statistical test results for RDSEED as the weak input source to the strong seeded randomness Circulant extractor. The seed is randomness generated from the NIST Randomness Beacon. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parenthesis.
@@
Table 20: Statistical test results for IDQ Quantis as the weak input source to the strong seeded randomness Circulant extractor. The seed is randomness generated from the NIST Randomness Beacon. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parenthesis.
@@ E.1 Two-source extraction with a single RNG
In this subsection, we test the use of two strings from each RNG as the inputs to a strong 2-source extractor. For near-perfect randomness to be generated, the unique strings from the RNG must be independent - otherwise this violates some of the assumptions of this level. Due to the limitations of the Circulant strong two-source extractor, we are unable to perform this step for the LFSR, since the min-entropy lower bound derived in Table 4 is too low.
@@
Table 21: This table gives the sum of tests failed for $5\times 10$Gbit samples from each RNG, after strong 2-source extraction taking strings of randomness from the same RNG and assuming independence. Note: The 32-bit LFSR does not generate any output in this setting, since it’s min-entropy is too low. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parenthesis.
@@
Table 22: Statistical test results for RDSEED as the weak input source to the strong two-source randomness Circulant extractor. The seed is randomness generated from RDSEED. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parenthesis.
@@
Table 23: Statistical test results for IDQ Quantis as the weak input source to the strong two-source randomness Circulant extractor. The seed is randomness generated from IDQ Quantis. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parenthesis.
@@ E.2 Two-source extraction using the NIST randomness beacon
Table 24: Statistical test results for the 32-bit LFSR as the weak input source to the strong two-source randomness Circulant extractor. The seed is randomness generated from the NIST Randomness Beacon. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parenthesis.
@@
Table 25: Statistical test results for RDSEED as the weak input source to the strong two-source randomness Circulant extractor. The seed is randomness generated from the NIST Randomness Beacon. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parenthesis.
@@
Table 26: Statistical test results for IDQ Quantis as the weak input source to the strong two-source randomness Circulant extractor. The seed is randomness generated from the NIST Randomness Beacon. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parenthesis.
@@ F.1 Protocol for physical randomness extraction
For this physical randomness extraction, we roughly follow the protocol developed in [46], with some adaptions to improve the randomness generation speed. This produces a semi-device-independent protocol for randomness amplification using a remote quantum computer, based on Bell tests. Roughly speaking, a Bell test requires a device to be challenged with inputs and then, based on the observed input-output statistics, a certain amount of entropy can be certified in the outputs. For a good description of Bell tests, see “Non-local games” in [52].
The adapted protocol we use is constructed as follows:
@@
    During each of the $n$ rounds, prepare a circuit that generates the GHZ state $\frac{1}{\sqrt{2}}(\left|000\right\rangle+i\left|111\right\rangle)$ and measure each qubit with a local $X$ or $Y$ measurement decided by the inputs at that round, that select from the set of measurements $\{(X,X,X),(X,Y,Y),(Y,X,Y),(Y,Y,X)\}$. Labelling local $X$ and $Y$ measurements as 00 and $1$ respectively allows us to write each measurement setting in the set as $(x_{i},y_{i},x_{i}\oplus y_{i})$ where subscript $i$ denotes the $i$-th round and $x_{i},y_{i}$ are input bits selected using the initial RNG. See Section 6.3 Implementations of Mermin inequality violations on quantum computers in [46].
@@
    Run the circuit of round $i\in 1,2,...,n$, recording the measurement settings $x_{i},y_{i},x_{i}\oplus y_{i}$ and measurement outcomes $a_{i},b_{i},c_{i}$ of that round.
@@
    Evaluate the Mermin inequality [64] value $\mathsf{M_{obs}}$, where
@@
    from the observed probability distribution, where $E_{x,y,x\oplus y}$ denotes the correlator for measurements $(x,y,x\oplus y)$, defined by:
@@
    Reduce $\mathsf{M_{obs}}$ to account for finite statistics using the Höeffding Inequality, using the relationship between $\mathsf{M_{obs}}$ and the ‘losing probability’, found at the beginning of Appendix A.2 of [46]. Let $\epsilon_{\rm est}$ be the estimation error and $\mathsf{M}$ be the true (asymptotic) value of the Mermin inequality for some I.I.D quantum device, then, we find $\mathsf{M_{adj}}$, such that $\Pr(\mathsf{M_{adj}}>\mathsf{M})\leq\epsilon_{\rm est}$ by defining
@@
    Based on the adjusted value $\mathsf{M_{adj}}$, evaluate the amount of min-entropy in the measurement outcomes of the quantum device. This is performed using the analytic expression in [65], which applies to 2 output bits per round. For details, see Section 4.3 Quantum devices, Bell tests, and guessing probabilities [46] and note that, the relationship between guessing probability and min-entropy can be found in Appendix A.3, Equation (28) of [46].
@@
    Take 2 of the 3 output bits (e.g. $\mathbf{a},\mathbf{b}$, discard $\mathbf{c}$) for randomness extraction.
@@
    Perform strong two-source randomness extraction using the quantum computer outputs $(\mathbf{a},\mathbf{b})$ with a fresh string of randomness from an RNG, if the sum of min-entropy’s of each bit string is high enough for extraction. The output is a near-perfect bit string, which we call the seed.
@@
    Repeatedly perform strong seeded extraction using the generated seed and fresh strings from the initial RNG, concatenating the output following the same logic as level 3: two-source extraction.
For the extractor implementation, we again use the ${\sf Circulant}$ extractor, and steps (8) and (9) can be viewed as analagous to level 3, where the source $Y$ is instead generated by the described quantum process. We used the $\mathsf{H}$1-1 Quantinuum ion-trap quantum computer as our quantum device, executing $1.8\times 10^{6}$ circuits to obtain a weakly random seed size of $3.6\times 10^{6}$ bits. This experiment took approximately 33.5 hours of quantum computing time. We obtain $\mathsf{M_{obs}}=3.83\rightarrow\mathsf{M_{adj}}\approx 3.75\rightarrow\alpha_{\mathsf{Q}}\geq 0.518$, where $\alpha_{\mathsf{Q}}$ is the min-entropy rate of the quantum computer outputs $(\mathbf{a},\mathbf{b})$. Note: The min-entropy of the LSFR was too low to perform physical extraction, as we need $\alpha_{\mathsf{RNG}}>1-\alpha_{\mathsf{Q}}$.

### v2

@@ 6 Discussion
Our statistical testing environment, $\mathsf{STE}$, software, documentation, and build file can be found at https://github.com/CQCL/random_test (accessed on 28 November 2024), and the randomness extractor software library Cryptomite can be found at https://github.com/CQCL/cryptomite (accessed on 28 November 2024) and in Foreman et al. (2024). These tools may be independently interesting, making both statistical testing and randomness extraction easy to use and openly accessible.
@@
A number of interesting future directions arise. It would be interesting to perform the statistical testing of other RNGs with our test environment to analyse how they perform when tested beyond what is required by standardisation bodies. Similarly, it would be interesting to include different post-processing methods than the ones that we have presented. One could use, for example, vetted conditioning components from NIST Turan et al. (2018) and compare their results to the ones obtained using information-theoretic randomness extractors.
We could have moved even further in our numerical testing but, because numerical tests consume substantial computational resources, decided to omit certain test suites from our analysis, including SPRNG Mascagni and Srinivasan (2000) and Crypt-X Gustafson et al. (1994). Moreover, we were recently made aware of the numerical tests BitReps Jamie Pont and RaBiGeTe Piras , which are also not included in $\mathsf{STE}$. It would be interesting to include these in the analysis to obtain an even more intense statistical testing environment.
[…]
Conceptualisation, C.F. and F.J.C.; methodology, C.F. and F.J.C.; software, R.Y.; validation, C.F., R.Y. and F.J.C.; formal analysis, C.F.; investigation, C.F. and F.J.C.; resources, F.J.C.; data curation, C.F.; writing—original draft preparation, C.F. and F.J.C.; writing—review and editing, C.F., R.Y. and F.J.C.; visualisation, C.F., R.Y. and F.J.C.; supervision, F.J.C.; project administration, C.F. and F.J.C. All authors have read and agreed to the published version of the manuscript.
[…]
This research received no external funding.
[…]
Not applicable.
[…]
The data presented in this study are openly available in https://github.com/CQCL/random_test (accessed on 28 November 2024).
Acknowledgements.
[…]
@@ A.1 Linear Feedback Shift Register (LFSR)
The LFSR is a class of pseudo-RNGs that is commonly used in applications due to its speed and ease of implementation in both software and hardware, e.g., Datta et al. (2017); Sahithi et al. (2012). Notably, LFSRs are used in cryptography, including in hashing and authentication Krawczyk (1994) and stream ciphers Ekdahl (2003). For this work, we implement the maximal period LFSR found in Panda et al. (2012).
This LFSR generates pseudo-randomness as follows. Let $s=b_{1},\ldots,b_{32}$ denote the initial 32-bit state, where $b_{i}$ denotes the $i=1,\ldots,32$th bit.
@@
    Initialise the LFSR with the 32-bit initial state $s=b_{1},b_{2},...,b_{31},b_{32}$.
@@
    Repeat steps (2–5) until the desired amount of bits has been generated.
@@ A.2 Intel RDSEED
Intel manufactures a hardware RNG based on thermal noise, which is present in their computer processing units. This true-RNG is constructed as follows, although a more in-depth description can be found in Jun and Kocher (1999); Hamburg et al. (2012).
@@
    Initial weak randomness is generated from an entropy source. This source is a self-clocking circuit designed such that, when the clock is running, the circuit enters a meta-stable state, which then resolves to one of two possible states—determined randomly by thermal noise. The state in which the circuit resolves is the random bit output from the entropy source. This self-clocking occurs irregularly at around 3 GHz.
@@
    Health and swellness checks are performed, which are very simple statistical tests, with the goal of detecting critical failures in the entropy source.
@@
    Cryptographic processing of the randomness with AES is performed.
A user requests randomness from the true-RNG via the RDSEED instruction. Using Intel’s true-RNG, we cannot directly output raw randomness from the entropy source; the best available option is RDSEED, which includes some level of post-processing (as described above). Independent analyses of the Intel true-RNG’s quality, including Hamburg et al. (2012); Shrimpton and Terashima (2015), report a min-entropy rate for RDSEED of around 0.65, closely matching our result (see Section 4).
@@ A.3 IDQ Quantis QRNG
The IDQ Quantis (USB) is a QRNG based on photons hitting a 50:50 beam splitter and being detected in position 0 (reflected) and 1 (transmitted). In principle, if all components are accurately modelled and the device is shielded from any outside influence, the output is perfectly random numbers due to the laws of quantum mechanics. We refer the reader to the IDQ Quantis QRNG brochure for further details of its construction Quantique (2004).
@@ B.1 Full Results: Statistical Testing
Table 9: The number of failed tests for the raw output from the 32-bit LFSR. Note that only 127/920 PractRand tests were run due to the numerous failings in the $2^{25}$ byte case. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parentheses.
@@
[…]
Table 10: The number of failed tests for the raw output from RDSEED. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parentheses.
@@
[…]
Table 11: The number of failed tests for the raw output from IDQ Quantis. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parentheses.
@@
[…]
Table 12: Cont.
RNG
NIST
Diehard
ENT
SmallCrush
Alphabit
Rabbit
PractRand
@@
B.2 Full Results: Min-Entropy Estimators
[…]
@@ B.2 Full Results: Min-Entropy Estimators
NIST Min-Entropy Estimator (/byte)
NIST Min-Entropy Estimator (/bit)
@@
[…]
Table 14: Observed NIST min-entropy estimators for RDSEED raw output.
@@
NIST Min-Entropy Estimator (/byte)
NIST Min-Entropy Estimator (/bit)
@@
[…]
@@
NIST Min-Entropy Estimator (/byte)
NIST Min-Entropy Estimator (/bit)
@@
[…]
Table 16: Cont.
RNG
NIST Min-Entropy Estimator (/byte)
NIST Min-Entropy Estimator (/bit)
@@ B.2.1 Deriving a Min-Entropy Lower Bound
In this subsection, we derive a min-entropy lower bound from the observed NIST min-entropy estimators in Table 4. We use subscripts to index a single min-entropy estimator test and superscript to index the RNG which the variable refers to. Let $\mathsf{est}_{i}$ denote the $i$th observed NIST min-entropy estimate per bit for a test, $i=1,\ldots,10$, and $\overline{\mathsf{est}}$ the average estimate per bit. The sample standard deviation $\sigma$ (using Bessel’s correction) is given by
@@
We compute the lower bound for the min-entropy rate, $\alpha^{\mathsf{RNG}}$, including a finite statistics correction term to lower-bound the true estimated min-entropy rate of each RNG with a high probability. Specifically, we desire
@@
where $\mathrm{\overline{est}^{\mathsf{RNG}}}$ is the average NIST min-entropy estimator for the RNG and $\sigma^{\mathsf{RNG}}$ is the observed sample standard deviation that satisfies Equation 18, giving $\epsilon_{\mathsf{est}}\approx 2^{-39}$. Here, we have made the assumption that the NIST min-entropy estimator results are normally distributed (which we believe is reasonable due to each test sample being generated a significant time apart) and used the standard probability density function for normally distributed variables.
@@ C.1 Full Results
[…]
Table 17: The number of failed tests for the output of the 32-bit LFSR after post-processing with the Von Neumann extractor. Note that only 127/920 PractRand tests were run due to the numerous failings in the $2^{25}$ byte case. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parentheses.
@@
[…]
Table 18: Cont.
RNG
NIST
Diehard
ENT
SmallCrush
Alphabit
Rabbit
PractRand
@@
[…]
Table 19: The number of failed tests for the output of RDSEED after post-processing with the Von Neumann extractor. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parentheses.
@@
[…]
Table 20: The number of failed tests for the output of IDQ Quantis after post-processing with the Von Neumann extractor. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parentheses.
@@ D.1 Full Results
Table 21: Statistical test results for the 32-bit LFSR as the weak input source to the strong seeded randomness Circulant extractor. The seed is randomness generated from the NIST Randomness Beacon. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parentheses.
@@
Table 22: Statistical test results for RDSEED as the weak input source to the strong seeded randomness Circulant extractor. The seed is randomness generated from the NIST Randomness Beacon. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parentheses.
@@
Table 23: Statistical test results for IDQ Quantis as the weak input source to the strong seeded randomness Circulant extractor. The seed is randomness generated from the NIST Randomness Beacon. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parentheses.
@@ E.1 Two-Source Extraction with a Single RNG
In this subsection, we test the use of two strings from each RNG as the inputs to a strong two-source extractor. For near-perfect randomness to be generated, the unique strings from the RNG must be independent; otherwise, this violates some of the assumptions of this level. Due to the limitations of the Circulant strong two-source extractor, we are unable to perform this step for the LFSR, since the min-entropy lower bound derived in Table 4 is too low.
[…]
Table 24: This table gives the sum of statistical tests failed for $5\times 10$Gbit samples from each RNG, after strong 2-source extraction taking strings of randomness from the same RNG and assuming independence. Note: The 32-bit LFSR does not generate any output in this setting, since its min-entropy is too low. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parentheses.
@@
[…]
Table 25: Statistical test results for RDSEED as the weak input source to the strong two-source randomness Circulant extractor. The seed is randomness generated from RDSEED. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parentheses.
@@
Table 26: Statistical test results for IDQ Quantis as the weak input source to the strong two-source randomness Circulant extractor. The seed is randomness generated from IDQ Quantis. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parentheses.
@@ E.2 Two-Source Extraction Using the NIST Randomness Beacon
Table 27: Statistical test results for the 32-bit LFSR as the weak input source to the strong two-source randomness Circulant extractor. The seed is randomness generated from the NIST Randomness Beacon. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parentheses.
@@
[…]
Table 28: Statistical test results for RDSEED as the weak input source to the strong two-source randomness Circulant extractor. The seed is randomness generated from the NIST Randomness Beacon. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parentheses.
@@
[…]
Table 29: Cont.
RNG
NIST
Diehard
ENT
SmallCrush
Alphabit
Rabbit
PractRand
@@
[…]
Table 30: Statistical test results for IDQ Quantis as the weak input source to the strong two-source randomness Circulant extractor. The seed is randomness generated from the NIST Randomness Beacon. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parentheses.
@@ F.1 Protocol For Physical Randomness Extraction
For this physical randomness extraction, we roughly follow the protocol developed in Foreman et al. (2023), with some adaptations to improve the randomness generation speed. (Other protocols could be used and adapted—for example Kessler and Arnon-Friedman (2020)—but additional analysis would need to be performed. Some other protocols allow for fewer assumptions than ours in Foreman et al. (2023),e.g., Chung et al. (2014); Ramanathan (2023), but these require significant classical computation so cannot be implemented on the current hardware). This produces a semi-device-independent protocol for randomness amplification using a remote quantum computer, based on Bell tests. Roughly speaking, a Bell test requires a device to be challenged with inputs and then, based on the observed input–output statistics, a certain amount of entropy can be certified in the outputs. For a good description of Bell tests, see “Non-local games” in Brunner et al. (2014).
The adapted protocol that we use is constructed as follows.
@@
[…]
@@
    Run the circuit of round $i\in 1,2,...,n$, recording the measurement settings $x_{i},y_{i},x_{i}\oplus y_{i}$ and measurement outcomes $a_{i},b_{i},c_{i}$ of this round.
@@
    Evaluate the Mermin inequality Mermin (1990) value $\mathsf{M_{obs}}$, where
@@
    from the observed probability distribution, where $E_{x,y,x\oplus y}$ denotes the correlator for measurements $(x,y,x\oplus y)$, defined by
@@
    Reduce $\mathsf{M_{obs}}$ to account for finite statistics using the Höeffding inequality, using the relationship between $\mathsf{M_{obs}}$ and the ‘losing probability’, found at the beginning of Appendix A.2 of Foreman et al. (2023). Let $\epsilon_{\rm est}$ be the estimation error and $\mathsf{M}$ be the true (asymptotic) value of the Mermin inequality for some I.I.D. quantum device; then, we find $\mathsf{M_{adj}}$ such that $\Pr(\mathsf{M_{adj}}>\mathsf{M})\leq\epsilon_{\rm est}$ by defining
@@
    Based on the adjusted value $\mathsf{M_{adj}}$, evaluate the amount of min-entropy in the measurement outcomes of the quantum device. This is performed using the analytic expression in Woodhead et al. (2018), which applies to two output bits per round. For details, see Section 4.3 Quantum devices, Bell tests, and guessing probabilities Foreman et al. (2023). Note that the relationship between the guessing probability and min-entropy can be found in Appendix A.3, Equation (28) of Foreman et al. (2023).
@@
    Take two of the three output bits (e.g., $\mathbf{a},\mathbf{b}$, discard $\mathbf{c}$) for randomness extraction.
@@
    Perform strong two-source randomness extraction using the quantum computer outputs $(\mathbf{a},\mathbf{b})$ with a fresh string of randomness from an RNG, if the sum of the min-entropies of each bit string is high enough for extraction. The output is a near-perfect bit string, which we call the seed.
@@
    Repeatedly perform strong seeded extraction using the generated seed and fresh strings from the initial RNG, concatenating the output following the same logic as level 3, i.e., two-source extraction.
For the extractor implementation, we again use the ${\sf Circulant}$ extractor, and steps (8) and (9) can be viewed as analogous to level 3, where the source $Y$ is instead generated by the described quantum process. We used the $\mathsf{H}$1-1 Quantinuum ion-trap quantum computer as our quantum device, executing $1.8\times 10^{6}$ circuits to obtain a weakly random seed of $3.6\times 10^{6}$ bits. This experiment took approximately 33.5 h of quantum computing time. We obtain $\mathsf{M_{obs}}=3.83\rightarrow\mathsf{M_{adj}}\approx 3.75\rightarrow\alpha_{\mathsf{Q}}\geq 0.518$, where $\alpha_{\mathsf{Q}}$ is the min-entropy rate of the quantum computer outputs $(\mathbf{a},\mathbf{b})$. We note that the min-entropy of the LSFR was too low to perform physical extraction, as a necessary condition is $\alpha_{\mathsf{RNG}}>1-\alpha_{\mathsf{Q}}$.